OpenStudy (anonymous):

Sue and Mary each have a $7000 loan with a 5% interest rate. Sue’s monthly payment is $161 with a loan term of 4 years, and Mary’s monthly payment is $113 with a loan term of 6 years. How much more in interest will Mary pay? A. $161 B. $408 C. $728 D. $1136

OpenStudy (amistre64):

this is the same question, different values

OpenStudy (amistre64):

read it too fast

OpenStudy (anonymous):

yeah this one is more complicated

OpenStudy (asad786):

113 times 6 =

OpenStudy (amistre64):

same concept tho As = 7000 + Si Bm = 7000 + Mi diffence between Mi and Si, given that A and B are the respective number of payments made and s and m are the respective payment amounts

OpenStudy (asad786):

161times 4 =

OpenStudy (anonymous):

644

OpenStudy (asad786):

the answer two both of those find the difference

OpenStudy (amistre64):

Si = As - 7000 -Mi = -Bm + 7000 ---------------- diff = As - Bm difference in payment still

OpenStudy (anonymous):

so 678-644?

OpenStudy (amistre64):

um, 4 is terms of years, not months ... we re making monthly payments

OpenStudy (anonymous):

oh so if there is 4 years there will be 48 months

OpenStudy (amistre64):

4(12)(161) - 6(12)(113)

OpenStudy (anonymous):

so -408

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

and that would be answer b right?

OpenStudy (amistre64):

i would agree

OpenStudy (anonymous):

Can you help me with another?

OpenStudy (anonymous):

What is the total cost of an easy-access loan of $1700 borrowed for 30 days with a 29% monthly fee?